OpenStudy (anonymous):

y varies indirectly as twice x. When x = 1, y = 3. Find y when x = 4

OpenStudy (anonymous):

y=2x+1 so x=4 means y=9

myininaya (myininaya):

\[y=\frac{k}{2 x}\] indirectly think division now use (1,3) to find k

OpenStudy (anonymous):

my answer choices are in fractions

OpenStudy (anonymous):

1/2, 3/4, 2/3, 1/4

OpenStudy (anonymous):

so Myininaya is right k=6 so the answer is 6/8=3/4

OpenStudy (anonymous):

thank you so muchh man

myininaya (myininaya):

woman*

myininaya (myininaya):

thank you so much woman*

myininaya (myininaya):

:) lol

OpenStudy (anonymous):

lol woman

OpenStudy (anonymous):

im sorry ms

OpenStudy (anonymous):

lol